OpenStudy (anonymous):

What is the equation of the line that passes through (–2, –3) and is perpendicular to 2x – 3y = 6?

OpenStudy (shoba):

first 2x - 3y = 6 3y = 2x - 6 y = 2/3x - 2 y=mx + c from this equation the slope ism= 2/3 and the intercept is c = -2 next find the equation of the line that parallel to y = 2/3x - 2 which pass through the point (-2,-3) can u solve it now?

OpenStudy (shoba):

you need to use point-slope formula... \[y - y _{1} = m(x-x _{1}) m=is the slope and (x _{1},y _{1}) is the given point\]

OpenStudy (shoba):

m1m2 = -1 (if it is perpendicular) m1 = 2/3 m2 = - 3/2 (new) x1,y1 = (-2,-3) so y + 3 = -3/2 (x+2) 2y + 6 = -3x -6 2y = -3x - 12 y = -3/2 x -6